# Reassessing Noise Augmentation Methods

### v3

@@ 3.2 Dataset
  • Reverberation: Applied to 10% of the files to simulate echo effects.
  • Background Noise: Random samples from the Freesound portion of the MUSAN corpus, which include 843 recordings of music, speech, and background noises [7]. The recordings are incorporated into the clean signal at random signal-to-noise ratios.
  • Speed Variations: The audio signal is re-sampled at three different speeds: reduced by 5%, original rate, and increased by 5%.
These augmentation techniques help simulate real-world acoustic environments and assess the robustness of ASR systems under varying conditions. Figure 2 illustrates the spectrum of a benign audio clip and its corresponding noisy counterpart.
@@ 3.3 Adversarial Attacks
In our experiments, we utilize the C&W attack, a prominent adversarial technique initially designed to target the DeepSpeech ASR system. The implementation by [31] for SpeechBrain enables us to evaluate the robustness of different SpeechBrain-based models. We generate 300 adversarial examples per architecture per model using the C&W attack.
For the C&W attack, we base our implementation on the robust_speech framework [31], adhering to the parameters defined therein for consistency and comparability, with the maximum number of iterations set to 3000. This allow us to systematically assess the vulnerability of SpeechBrain models to sophisticated adversarial threats.
@@ 3.4 Statistical Tests
To assess whether adversarial samples generated for two different training regimes have the same underlying distribution, we apply the Kolmogorov-Smirnov (KS) test. We use a significance level of $\alpha=0.05$. For $p\leq 0.05$, any KS statistic above $0.1$ indicates a statistically significant difference.
@@ 4.1 Evaluating Noise Robustness
In our initial experiment, we assess the effectiveness of noise-augmented training on enhancing noise robustness in SpeechBrain models. For each architecture, we verify whether this training method improves the models’ performance against natural acoustic perturbations, as described in subsection 3.2. The evaluation involves two metrics:
    $\text{WER}=100\cdot\frac{S+D+I}{N}\enspace,$ (1)
@@
    where $N_{S}$ is the number of audio clips with no transcription error compared to the reference text, out of a total of $N_{T}$ examples. For benign samples, the ground truth serves as the reference, whereas for C&W adversarial examples, we use the target transcription.
4.2 Evaluating Adversarial Robustness
We employ two distinct types of distortion measurements to quantify the imperceptibility of adversarial perturbations, offering insights into the audibility of the attacks. Let $x(t)$ denote the original continuous audio file and $X(k)$ denote the discretized version:
[…]
    $\text{dB}(x)=20\cdot\max_{t}\log_{10}x(t)\text{.}$
    $\text{dB}_{\text{x}}(\delta)=\text{dB}(x)-\text{dB}(\delta)\text{,}$
    assuming equal lengths of $x$ and $\delta$.
[…]
    $\text{SNR${}_{\text{Seg}}$}=\frac{10}{M}\cdot\sum_{m=0}^{M-1}\log_{10}\frac{\sum_{k=mF}^{mF+F-1}X(k)^{2}}{\sum_{k=mF}^{mF+F-1}\delta(k)^{2}}\text{,}\enspace$
[…]
@@ 5 Results
In the following we present and discuss the results of our experiments. A demo is available at https://matiuste.github.io/SPSC_24/.
@@
Wav2vec 1 85.67% 3.11% 81.33% 5.39% 2.28%
2 87.00% 2.84% 83.00% 5.06% 2.22%
3 86.33% 2.95% 82.33% 5.12% 2.17%
[…]
2 71.33% 8.06% 55.67% 20.02% 11.96%
3 71.00% 8.57% 65.33% 11.51% 2.95%
[…]
2 71.33% 8.18% 54.67% 20.02% 11.85%
3 66.67% 10.01% 56.67% 14.35% 4.34%
[…]
2 66.67% 9.29% 53.00% 19.47% 10.18%
3 68.00% 8.51% 53.67% 18.85% 10.34%
@@
Wav2vec 1 100.0% 0.00 155 35.41 18.54
2 99.6% 0.05 343 29.59 14.08
3 100.0% 0.00 158 35.40 18.28
@@
2 86.0% 7.16 323 27.46 9.47
3 74.0% 14.89 479 23.92 6.69
[…]
2 99.6% 0.05 113 41.13 18.84
3 96.0% 1.53 186 37.65 15.93
[…]
2 100.0% 0.00 18 54.52 27.23
3 100.0% 0.00 20 54.90 27.59
@@
Table 2: Results for C&W targeted attack. All values are averaged over 300 adversarial examples. WER and success rate are measured w.r.t. the target adversarial transcription. The arrows indicate the direction for models that are more robust against adversarial examples, e.g., models with a lower success rate are considered more robust against adversarial examples. The bold numbers indicate the more robust model with respect to the chosen parameter. Underlined numbers show a statistical significant difference between training regime 1 and the considered training regime according to the KS test.
5.1 Evaluating Noise Robustness
In evaluating noise robustness, the reference text for WER and success rate calculations is the ground-truth transcript of the clean test sample. We aim for ASR models to exhibit low WER and high success rate, indicating accurate recognition of the ground-truth transcription. This step is essential to validate the effectiveness of our noise-augmented training regimen before examining the relationship between noise and adversarial robustness.
Our experimental results corroborate that models subjected to noise-augmented training exhibit enhanced robustness when confronted with speech samples with environmental noise. For each model considered in Table 1, the success rate and WER for noisy data have improved with noise-augmented training (the models marked in blue or red). This is evidenced by a lower WER in comparison to their counterparts trained without any augmentation (marked in black).
This improvement is particularly pronounced for the seq2seq models, which build on minimal pre-trained components, while the enhancement is less apparent for the wav2vec and transformer models that rely more on pre-trained parts, especially for feature extraction. Interestingly, the wav2vec-based models maintain a high success rate even with noisy data, whereas for all other architectures, the success rate drops significantly. This could be attributed to the fact that the wav2vec models incorporate pre-trained models that were exposed to noisy data during training. These findings indicate that noise-augmented training indeed enhances noise robustness, although the extent of improvement varies depending on the architecture.
In the next experiment, we evaluate if it also enhances adversarial robustness.
5.2 Evaluating Adversarial Robustness
In the subsequent experiment, we focus on assessing adversarial robustness. For every C&W adversarial example we gather data similar to [46]. To gauge the severity of performance degradation, we compare the success rate of adversarial attacks against our models (see Eq. 2) and analyze the WER (see Eq. 1). Additionally, we evaluate the computational complexity of generating adversarial examples by recording the first hit, defined as the number of iterations required to craft the first successful adversarial example.
Table 2 encapsulates the success rate, computational effort, and perceptibility metrics across various architectures and models for the targeted C&W attack. Our findings reveal that for the seq2seq models trained from scratch, those that underwent noise-augmented training demonstrate a slightly reduced adversarial success rate, a higher WER, or more perceptible noise, indicating improved robustness to adversarial attacks. Again, the effects on improved adversarial robustness are most apparent for the seq2seq-based models. Similarly to the noise robustness, only minor improvements can be seen for the wav2vec- and transformer-based models. Interestingly, we did not find a clear connection between the model’s WER on the benign data (see Table 1, fourth column) and the adversarial success rate (see Table 2, third column).
@@ 5.2 Evaluating Adversarial Robustness
Our findings indicate that incorporating speed or noise augmentation into the training process enhances the adversarial robustness of most models when confronted with targeted white-box attacks. In cases where data augmentation does not lead to higher robustness compared to the base model, it also does not result in a severe robustness reduction. This underlines that noise augmentation should be used as a standard routine in the training of neural network-based ASR systems.
6 Discussion and Conclusion
Our comprehensive study has shed light on the efficacy of noise augmentation as a strategy for improving the adversarial robustness of automatic speech recognition (ASR) systems. By conducting a systematic comparative analysis of four state-of-the-art ASR architectures, each trained with different augmentation techniques and subjected to targeted white-box attacks, we have provided empirical evidence supporting the benefits of noise augmentation.

### v4

@@ 3.3 Data Augmentation
  • 1.
    Reverberation: To $10\%$ of the files, we apply reverberation to simulate echo effects.
  • 2.
    Background noise: Random samples from the Freesound portion of the MUSAN corpus, which include $843$ recordings of music, speech, and background noises (Ko et al., 2017). The recordings are incorporated into the clean signal at random signal-to-noise ratios.
  • 3.
    Speed variations: The audio signal is re-sampled at three different speeds: reduced by $5\%$, original rate, and increased by $5\%$.
Figure 4 illustrates the spectrum of a benign audio signal and its corresponding noisy counterpart. See A for the code snippet that shows how the corruptions are employed.
(a) Original audio signal. Transcription: ”to their sorrow they were soon undeceived”.
(b) Noisy original audio signal. Transcription: ”to their sorrow they were assumes under seats”.
(c) Noise signal. Difference between (a) and (b).
Figure 4: The spectrogram of the original audio signal (a) is compared to the spectrogram of its corresponding noisy example, that in this case is constructed by adding background noise to the original audio signal (b) and the spectrogram of the noise (c), which is the difference between the original and its noisy counterpart.
@@ 3.4 Adversarial Attacks
In our experiments, we utilize the C&W white-box attack, a prominent adversarial technique initially designed to attack the DeepSpeech ASR system, see Section 2.1. We generate 300 adversarial examples per architecture and per model. In addition, we investigate two black-box attacks: the Alzantot and the Kenansville attack, for each of which we create 100 adversarial examples per architecture and per model.
For C&W and Alzantot attacks, we base our implementation on the robust_speech framework (Olivier and Raj, 2022), adhering to the parameters defined therein for consistency and comparability, with the maximum number of iterations set to $3000$. These methodologies allow us to systematically assess the vulnerability of SpeechBrain models to sophisticated adversarial threats. For the Kenansville attack, we use its online implementation [^4] Kenansville attack code, see https://github.com/kwarren9413/kenansville_attack. and run 15 iterations.
3.5 Evaluation metrics
To evaluate the ASR performance on benign, noise-augmented, or adversarial files, we use the following two metrics:
  • 1.
    $\text{WER}=100\cdot\frac{S+D+I}{N}\kern 5.0pt,$ (1)
  • 2.
    $\text{Succ.\ rate}=100\cdot\frac{N_{S}}{N_{T}}\kern 5.0pt,$ (2)
    where $N_{S}$ is the number of audio clips with no transcription error compared to the reference text, out of a total of $N_{T}$ examples. For benign samples and black-box adversarial attacks, the ground truth serves as the reference, whereas for C&W adversarial examples, we use the target transcription. For adversarial success rate, we distinguish between two cases: for targeted attacks, the reference is the target sentence; for untargeted attacks, any transcription differing from the ground truth sentence counts as an adversarial success.
In order to quantify the distortion in adversarial audio files, we use the following metrics. Let $x(t)$ denote the original continuous audio file and $X(k)$ denote the discretized version:
  • 1.
[…]
    $\text{dB}(x)=20\cdot\max_{t}\log_{10}x(t)\kern 5.0pt{.}$
    $\text{dB}_{\text{x}}(\delta)=\text{dB}(x)-\text{dB}(\delta)\kern 5.0pt{,}$
    assuming equal lengths of $x$ and $\delta$.
  • 2.
    SI-SDR: The Scale-Invariant Signal-to-Distortion Ratio improves upon the traditional non-segmental Signal-to-Noise Ratio by considering the scale-invariance of audio signals (Roux et al., 2019). It measures, in decibels, the ratio of the energies of the reference sources to the distortion signal, while accounting for any scaling differences between the two signals. The SI-SDR is calculated as
    $\text{SI-SDR}=10\cdot\log_{10}\frac{\sum_{k}(\alpha\cdot X(k))^{2}}{\sum_{t}(\alpha\cdot X(k)-\hat{X}(k))^{2}}\kern 5.0pt{,}$
    where $X$ represents the benign example, and $\hat{X}$ is its adversarial counterpart and the scaling factor $\alpha$ is given by
    $\alpha=\frac{\sum_{k}X(k)\cdot\hat{X}(k)}{\sum_{k}X(k)^{2}}\kern 5.0pt{.}$
    The calculation assumes that $X$ and $\hat{X}$ are of equal length.
  • 3.
[…]
    $\text{SNR${}_{\text{Seg}}$}=\frac{10}{M}\cdot\sum_{m=0}^{M-1}\log_{10}\frac{\sum_{k=mF}^{mF+F-1}X(k)^{2}}{\sum_{k=mF}^{mF+F-1}\delta(k)^{2}}\text{,}\kern 5.0pt$
@@ 3.6 Statistical Tests
To assess whether adversarial samples generated from two different training regimes share the same underlying distribution, we perform the Kolmogorov-Smirnov (KS) test. This non-parametric test compares the distributions of two samples and tests the null hypothesis that they are drawn from the same distribution. The hypothesis for the KS test is as follows:
  • 1.
    Null hypothesis: The two distributions are identical.
  • 2.
    Alternative hypothesis: The two distributions are different.
A significance level of $\alpha=0.05$ is used for the test. If the p-value is less than or equal to $0.05$ ($p\leq 0.05$), and the Kolmogorov-Smirnov statistic exceeds $0.1$, it indicates that the distributions are significantly different. On the other hand, if the p-value is greater than 0.05 ($p>0.05$), we fail to reject the null hypothesis, implying that there is no statistically significant difference between the two distributions. Therefore, if the KS test reveals a statistically significant difference, it suggests that the adversarial samples from the two regimes differ significantly. This could imply a noticeable change in model robustness due to the differences in training regimes. For instance, if one regime shows significantly better adversarial performance, it may indicate that its training setup contributes positively to the model’s robustness.
@@ 4.1 Evaluating Noise Robustness
In our initial experiment, we assess the effectiveness of noise-augmented training on enhancing noise robustness in SpeechBrain models. For each architecture, we verify whether robustness training improves the models’ performance against natural acoustic perturbations. For this, we add background noise to every sample according to the guidelines specified in the “Background Noise” from subsection 3.3. For detailed information on the parameter setup, refer to B.
In evaluating noise robustness, the reference text for WER and success rate calculations is the ground-truth transcript of the clean test sample. When training ARS models, we aim for models that exhibit low WER and high benign success rate on benign and noise-augmented data, indicating accurate recognition of the ground-truth transcription.
4.2 Evaluating Adversarial Robustness
To assess adversarial robustness, we first generated adversarial examples for the three different attack types (C&W, Alzantot, and Kenansville) described above. To assess the strength of the adversarial examples, we compare the success rate of adversarial attacks against our models (see Eq. 2) and analyze the WER (see Eq. 1). For the untargeted black-box adversarial attacks, the ground truth serves as the reference; that is, the attack is considered stronger for a higher WER and lower success rate. For adversarial examples generated by the targeted white-box attack C&W, we use the target transcription as ground truth. Thus, the lower the WER and the higher the success rate, the stronger the attack.
Additionally, for the C&W and Alzantot attacks, we evaluate the computational complexity of generating adversarial examples by recording the first hit, defined as the number of iterations required to craft the first successful adversarial example.
@@ 4.2 Evaluating Adversarial Robustness
Tracking this set of different metrics allows us to dissect the nuances of adversarial robustness, understanding the trade-offs between attack efficacy, computational expense, and perceptibility in the context of our SpeechBrain models.
@@ 5 Results
In the following, we present and discuss the results of our experiments. A demo is available at https://matiuste.github.io/AssesNoiseAugm/, where we also provide pre-trained models to support reproducibility. These models can be used to generate adversarial examples and evaluate the effectiveness of noise augmentation methods.
@@
CTC No augmentation 85.67% 3.11% 81.33% 5.39% 2.28%
Speed-augmented 87.00% 2.84% 83.00% 5.06% 2.22%
Fully augmented 86.33% 2.95% 82.33% 5.12% 2.17%
Seq2seq 1 No augmentation 70.33% 8.18% 54.33% 19.52% 11.35%
Speed-augmented 71.33% 8.06% 55.67% 20.02% 11.96%
Fully augmented 71.00% 8.57% 65.33% 11.51% 2.95%
Seq2seq 2 No augmentation 67.00% 9.23% 53.33% 21.08% 11.85%
Speed-augmented 71.33% 8.18% 54.67% 20.02% 11.85%
Fully augmented 66.67% 10.01% 56.67% 14.35% 4.34%
Transformer No augmentation 65.00% 9.90% 52.33% 19.91% 10.01%
Speed-augmented 66.67% 9.29% 53.00% 19.47% 10.18%
Fully augmented 68.00% 8.51% 53.67% 18.85% 10.34%
@@
5.1 Evaluating Noise Robustness
To evaluate noise robustness, WER and success rate are calculated using the ground-truth transcript of the clean test sample. Low WER and high success rate indicate accurate transcription, allowing us to assess the effectiveness of noise-augmented training before analyzing its impact on adversarial robustness.
Our experimental results, see Table 1, corroborate that models subjected to noise-augmented training exhibit enhanced robustness when confronted with speech samples with environmental noise. For each model considered, the success rate and WER for noisy data have improved with noise-augmented training (the most robust models are highlighted in bold). This is evidenced by a lower WER in comparison to their counterparts trained without any augmentation.
This improvement is particularly pronounced for the seq2seq models, which build on minimal pre-trained components, while the enhancement is less apparent for the transformer-based models (including CTC) that rely more on pre-trained parts, especially for feature extraction. Interestingly, the CTC-based models maintain a high success rate even with noisy data, whereas for all other architectures, the success rate drops significantly. This could be attributed to the fact that the CTC models incorporate pre-trained models that were exposed to noisy data during training. These findings indicate that noise-augmented training indeed enhances noise robustness, although the extent of improvement varies depending on the architecture.
@@ 5.1 Evaluating Noise Robustness
CTC No augmentation 100.0% 0.00 155 35.41 29.29 18.54
Speed-augmented 99.6% 0.05 343 29.59 25.30 14.08
Fully augmented 100.0% 0.00 158 35.40 29.04 18.28
Seq2seq 1 No augmentation 88.3% 5.53 454 27.21 20.29 9.34
Speed-augmented 86.0% 7.16 323 27.46 20.42 9.47
Fully augmented 74.0% 14.89 479 23.92 17.74 6.69
Seq2seq 2 No augmentation 98.3% 0.47 107 41.99 30.23 19.37
Speed-augmented 99.6% 0.05 113 41.13 29.65 18.84
Fully augmented 96.0% 1.53 186 37.65 26.88 15.93
Transformer No augmentation 100.0% 0.00 19 56.26 40.77 28.89
Speed-augmented 100.0% 0.00 18 54.52 39.10 27.23
Fully augmented 100.0% 0.00 20 54.90 39.47 27.59
Table 2: Results for C&W targeted attack. All values are averaged over 300 adversarial examples. WER and success rate are measured w.r.t. the target adversarial transcription. The arrows indicate the direction for models that are more robust against adversarial examples, e.g., models with a lower success rate are considered more robust against adversarial examples. The bold numbers indicate the more robust model with respect to the chosen parameter. Underlined numbers indicate a statistically significant difference based on the KS test between the no augmentation training regime and the evaluated augmented training regime.
5.2 Evaluating Adversarial Robustness
Table 2 encapsulates the success rate, computational effort, and perceptibility metrics across various architectures and models for the targeted C&W attack. In this setting, WER and success rate are computed using the target adversarial transcript, where higher WER and lower success rate indicate stronger adversarial robustness. For distortion metrics, lower values of $\text{dB}_{x}$, SI-SDR, and $\text{SNR}_{\text{seg}}$ suggest that greater noise is needed for the adversarial samples to succeed, reflecting better model resistance.
The results reveal that for the seq2seq models trained from scratch, those that underwent noise-augmented training demonstrate a slightly reduced adversarial success rate, a higher WER, or more perceptible noise, indicating improved robustness to adversarial attacks. As shown in Table 2, the underlined values indicate cases where distortion metrics differ statistically significantly between the baseline (no augmentation) and data augmentation regimes, indicating that data augmentation leads to noisier adversarial examples. While improvements in WER and success rate are modest, all augmentation regimes consistently enhance adversarial robustness by forcing adversarial examples to be noisier. This effect is most pronounced in seq2seq-based models, which benefit significantly from augmentation. Transformer-based models (including the CTC model) show more limited gains but still exhibit increased resistance to adversarial attacks, mirroring the trends observed in noise robustness. Interestingly, we did not find a clear connection between the model’s WER on the benign data (see Table 1, fourth column) and the adversarial success rate (see Table 2, third column).
Alzantot Adversarial Data
Model Augm. Succ. rate $\downarrow$ WER $\downarrow$ First hit $\uparrow$ SI-SDR $\downarrow$ $\text{SNR}_{\text{seg}}$ $\downarrow$
CTC No augmentation 58.00% 10.63% 77 20.75 12.25
Speed-augmented 61.00% 12.18% 91 18.95 8.13
Fully augmented 60.00% 10.94% 56 21.28 12.10
Seq2seq 1 No augmentation 96.00% 45.92% 42 18.61 8.15
Speed-augmented 95.00% 42.72% 37 19.40 8.72
Fully augmented 86.00% 29.31% 77 18.97 8.52
Seq2seq 2 No augmentation 96.00% 51.29% 37 17.86 7.05
Speed-augmented 99.00% 47.78% 37 18.66 8.05
Fully augmented 91.00% 33.13% 72 18.34 7.94
Transformer No augmentation 99.00% 39.63% 42 19.99 9.24
Speed-augmented 96.00% 38.18% 50 20.80 11.24
Fully augmented 99.00% 36.74% 34 20.86 11.12
Table 3: Results for the Alzantot untargeted attack. Values are averaged over 100 adversarial examples. For this untargeted attack, when calculating the adversarial success rate, any transcription but the ground truth transcription is considered a success. For simplicity, the WER is measured w.r.t. the ground truth transcription. It thus provides an insight of how different the resulting adversarial samples are from the ground truth. Again, the arrows indicate the direction for models that are more robust against adversarial examples. The bold numbers indicate the more robust model with respect to the chosen parameter. Underlined numbers indicate a statistically significant difference based on the KS test between the no augmentation training regime and the evaluated augmented training regime.
Kenansville Adversarial Data
@@ 5.2 Evaluating Adversarial Robustness
CTC No augmentation 80.00% 13.62% 12.90 14.94 6.58
Speed-augmented 99.00% 17.75% 10.13 12.58 4.56
Fully augmented 78.00% 13.83% 11.89 14.32 5.98
Seq2seq 1 No augmentation 92.00% 21.88% 19.28 21.39 12.16
Speed-augmented 96.00% 24.97% 18.69 20.72 11.46
Fully augmented 99.00% 22.81% 17.21 19.44 10.56
Seq2seq 2 No augmentation 97.00% 31.37% 18.38 20.80 11.67
Speed-augmented 94.00% 24.77% 19.11 21.49 12.32
Fully augmented 96.00% 26.83% 15.83 18.34 9.56
Transformer No augmentation 96.00% 24.46% 20.75 22.77 13.46
Speed-augmented 97.00% 22.81% 19.56 21.80 12.60
Fully augmented 96.00% 21.57% 20.22 22.32 13.02
@@
Table 4: Results for the Kenansville untargeted attack. All values are averaged over 100 adversarial examples. For this untargeted attack, when calculating the adversarial success rate, any transcription but the ground truth transcription is considered a success. For simplicity, the WER is measured w.r.t. the ground truth transcription. It thus provides an insight of how different the resulting adversarial samples are from the ground truth. The arrows indicate the direction for models that are more robust against adversarial examples. The bold numbers indicate the more robust model with respect to the chosen parameter. Underlined numbers indicate a statistically significant difference based on the KS test between the no augmentation training regime and the evaluated augmented training regime.
While from a privacy and security perspective, the most relevant attacks are targeted, we also explored the impact of noise-augmented training on the model robustness w.r.t. untargeted black-box attacks. In this context, WER and success rate are calculated using the ground-truth transcript, where lower WER and success rate reflect stronger adversarial robustness. While the success rate defines the attacks’ success, the WER is a general measure of transcription accuracy. For distortion metrics, lower values indicate that adversarial samples need more noise to succeed, demonstrating better model resilience. Here, the benefits of noise-augmented training are less definitive, as detailed in Tables 3 and 4. For the Alzantot attack, models trained with augmented data exhibit commendable performance in terms of success rate and WER, with the seq2seq augmented models displaying significant improvements, as highlighted by the underlined WER values in Table 3. These models demonstrate enhanced robustness, as confirmed by the Kolmogorov-Smirnov test, showing statistically significant differences compared to the no augmentation counterparts. While models trained without augmentation show a slight advantage in distortion metrics like segmental SNR, this difference is not statistically significant. For the Kenansville attack, differences in success rate and WER between augmented and non-augmented models are not statistically significant. However, models trained with augmented data consistently show lower distortion metrics, indicating improved robustness. The seq2seq models, in particular, demonstrate statistically significant improvements, as underlined in Table 4. Overall, augmented models are more resistant to adversarial attacks, requiring adversarial examples to be noisier in order to succeed.
@@
In summary, our findings indicate that incorporating speed or noise augmentation into the training process enhances the adversarial robustness of most models when confronted with targeted white-box attacks as well as untargeted black-box attacks. In cases where data augmentation does not lead to higher robustness compared to the base model, it also does not result in a severe robustness reduction. This underlines that noise augmentation should be used as a standard routine in the training of neural network-based ASR systems.
6 Discussion and Future Work